# A non-abelian version of Deligne’s Fixed Part Theorem

### v4

@@ 4.2. Admissible representations
In this subsection we assume that $K$ is algebraically closed of characteristic zero. A representation $\rho_{H}\colon H\to{\mathrm{GL}}(V)$ is called admissible if it is semi-simple and all simple constituents of $\rho_{H}$ have finite determinant. A representation $\rho\colon G\to{\mathrm{GL}}(V)$ is called admissible relative to $H$ (or an admissible extension of $\rho|_{H}$) if $\rho$ is admissible and all its simple constituents stay simple when restricted to $H$.
@@
The algebraic monodromy group $\mathrm{Mo}(\rho)^{0}$ of $\rho$ is semi-simple if and only if $\rho|_{G^{\prime}}$ is admissible for all subgroups of finite index $G^{\prime}\subset G$.
@@
Assume the algebraic monodromy group $\mathrm{Mo}(\rho)^{0}$ is semi-simple. Then for all finite index subgroups $G^{\prime}\subset G$, $\rho(G^{\prime})\subset\rho(G)$ has finite index, thus $\mathrm{Mo}(\rho|_{G^{\prime}})\subset\mathrm{Mo}(\rho)$ has finite index as well. This implies the equality $\mathrm{Mo}(\rho|_{G^{\prime}})^{0}=\mathrm{Mo}(\rho)^{0}$.
In particular $\mathrm{Mo}(\rho|_{G^{\prime}})^{0}$ is reductive, thus by Weyl’s theorem, $\rho|_{G^{\prime}}$ as a representation of the algebraic group ${\rm Mo}(\rho|_{G^{\prime}})^{0}$ is semi-simple, thus $\rho|_{G^{\prime}}$ itself is semi-simple. As ${\rm Mo}(\rho)^{0}=[{\rm Mo}(\rho)^{0},{\rm Mo}(\rho)^{0}]$, all rank one algebraic representations of ${\rm Mo}(\rho|_{G^{\prime}})$ are finite. Thus $\rho|_{G^{\prime}}$ is admissible.
Assume conversely that $\rho|_{G^{\prime}}$ is admissible for all subgroups of finite index $G^{\prime}\subset G$. Upon replacing $G$ by the preimage of ${\rm Mo}(\rho)^{0}(K)$, which is a finite index subgroup in $G$ (recall $K$ is algebraically closed), we may assume that $\mathrm{Mo}(\rho)$ is connected. As the unipotent radical $U$ of $\mathrm{Mo}(\rho)$ is normal, by Clifford’s theorem $\rho|_{U}$ is still semi-simple. So by Engel’s theorem, we conclude that $U$ is trivial. The radical $R$ of $\mathrm{Mo}(\rho)$ is then a central torus, so by Schur’s lemma it acts by a character $\lambda\colon R\to\mathbb{G}_{m}$ on each simple $G$-subrepresentation $V_{i}\subset V$. As the action on $\det(V_{i})$ is finite, a power of $\lambda$ is trivial, so $\lambda$ is trivial. As the action of $R$ on $V$ is faithful, $R$ is trivial and ${\rm Mo}(\rho)^{0}$ is semi-simple. ∎
@@
Finite direct sums of representations with semi-simple algebraic monodromy groups have semi-simple algebraic monodromy groups. Direct summands of representations with semi-simple algebraic monodromy group have a semi-simple algebraic monodromy group.
@@
By Lemma 4.2 it is sufficient to prove this for “semi-simple algebraic monodromy group” replaced by “admissible in restriction to all finite index subgroups”, for which it is trivial. ∎
@@
for all $h\in H$ and $g\in G$, and as $\rho|_{H}$ is simple, Schur’s lemma implies that $\tau(g)\in K^{\times}$ for all $g\in G$. In fact $\tau(g)\in\mu_{N}(K)$ for some $N>0$ as the determinants of $\rho$ and $\tilde{\rho}$ are finite. This implies that $\tau\colon G\to\mu_{N}(K)\subset{\mathrm{GL}}(V)$ is a homomorphism. Thus $\rho=\tilde{\rho}$ in restriction to the kernel $G^{\prime}\subset G$ of $\tau$, which is a finite index subgroup. ∎
@@ 4.3. Algebraically isomonodromic representations
If $\rho$ is algebraically isomonodromic relative to $H$, then $\mathrm{Mo}(\rho|_{H})\to\mathrm{Mo}(\rho)$ is essentially an isomorphism. In particular $\mathrm{End}_{\rho}(V)\xrightarrow{\sim}\mathrm{End}_{\rho|_{H}}(V)$ is essentially an isomorphism.
@@
Assume $K$ is algebraically closed. If $\rho\colon G\to{\mathrm{GL}}(V)$ is algebraically isomonodromic relative to $H$ and $\mathrm{Mo}(\rho)^{0}$ semi-simple, then $\rho$ is essentially admissible relative to $H$.
@@
Assume $K$ is algebraically closed. Let $\rho\colon G\to{\mathrm{GL}}(V)$ be a represenation such that $\mathrm{Mo}(\rho|_{H})$ is connected and semi-simple. Then $\rho$ is essentially admissible relative to $H$ if and only if $\rho$ is algebraically isomonodromic relative to $H$.
@@
For $\Rightarrow$” we first observe that by replacing $G$ by the preimage of $\mathrm{Mo}(\rho)^{0}$, which is a subgroup of finite index containing $H$, we may assume that $\mathrm{Mo}(\rho)$ is connected. The argument in the proof of Lemma 4.2 shows that $\mathrm{Mo}(\rho)$ is semi-simple. If the embedding $\mathrm{Mo}(\rho|_{H})\to\mathrm{Mo}(\rho)$ is not an isomorphism then there exists a non-trivial normal connected algebraic subgroup $W\subset\mathrm{Mo}(\rho)$ centralizing $\mathrm{Mo}(\rho|_{H})$ such that
@@
is surjective with finite kernel, see e.g. [Hum75, 14.2]. The $W$ action stabilizes every simple $\rho$-subrepresentation $V^{\prime}\subset V$ and commutes with the action of $H$ on $V^{\prime}$, which is simple. So by Schur’s lemma $W$ acts by a character on $V^{\prime}$, but all characters are trivial as $W$ is semi-simple. So $W$ has to be trivial as it acts faithfully on $V$. ∎
@@
Assume that $K$ is algebraically closed. Let $\rho_{H}\colon H\to{\mathrm{GL}}(V)$ be a representation with semi-simple algebraic monodromy group. Let $G^{\prime}\subset G$ be a subgroup of finite index, $H^{\prime}=H\cap G^{\prime}$. If $\mathrm{Mo}(\rho|_{H^{\prime}})$ is connected and $\rho|_{H^{\prime}}$ has an admissible extension $\rho^{\prime}\colon G^{\prime}\to{\mathrm{GL}}(V)$ relative to $H^{\prime}$, then $\rho$ essentially has an algebraically isomonodromic extension $\rho\colon G\to{\mathrm{GL}}(V)$ relative to $H$.
@@
After replacing $G^{\prime}$ by a subgroup of finite index we may assume that $G^{\prime}\subset G$ is normal, thus $H^{\prime}\subset H$ is normal as well. Replace $G$ by the preimage of $H/H^{\prime}\subset G/G^{\prime}$ via $G\to G/G^{\prime}$. Then by definition $[G:G^{\prime}]=[H:H^{\prime}]$. We set
@@
We show in the sequel that there is a direct summand of $\tilde{\rho}$ which extends $\rho_{H}$. First observe that $\tilde{\rho}$ is algebraically isomonodromic relative to $H$. Indeed, it suffices to check that $\tilde{\rho}|_{G^{\prime}}$ is algebraically isomonodromic relative to $H^{\prime}$. Noting that $\tilde{\rho}|_{H^{\prime}}$ is a direct sum of copies of $\rho_{H}|_{H^{\prime}}$, so $\mathrm{Mo}(\tilde{\rho}|_{H^{\prime}})$ is connected and semi-simple, by Proposition 4.10 we have to show that $\tilde{\rho}|_{G^{\prime}}$ is admissible relative to $H^{\prime}$. This holds as $\tilde{\rho}|_{G^{\prime}}$ it is a direct sum of $H$-conjugates of $\rho^{\prime}$ and $\rho^{\prime}$ is admissible relative to $H^{\prime}$. On the other hand,
@@
We clearly have c) $\Rightarrow$ b) $\Rightarrow$ a). We now reduce the proof of the implication a) $\Rightarrow$ c) to the case where $K$ is algebraically closed. Let $\overline{K}$ be an algebraic closure of $K$. Consider an algebraically isomonodromic representation $\rho\colon G\to{\mathrm{GL}}(V\otimes_{K}\overline{K})$ relative to $H$ with $\rho|_{H}$ stabilizing $V$ (thus defined over $K$). Then, as $G$ is finitely generated, there exists a finite Galois subextension $\tilde{K}\subset\overline{K}$ of $K$ such that $\rho$ stabilizes $V\otimes_{K}\tilde{K}$. The Galois group ${\rm Gal}(\tilde{K}/K)$ acts on $V\otimes_{K}\tilde{K}$ via its action on $\tilde{K}$, thus on ${\mathrm{GL}}(V\otimes_{K}\tilde{K})$. For $\gamma\in{\rm Gal}(\tilde{K}/K)$ we define $\gamma\cdot\rho$ by the formula $\gamma\cdot\rho(g)=\gamma\cdot(\rho(g))$. Then ${\rm Mo}(\gamma\cdot\rho)=\gamma\cdot{\rm Mo}(\rho)$ thus both $\gamma\cdot\rho$ and $\rho$ are algebraically isomonodromic extensions relative to $H$ of $\rho|_{H}=\gamma\cdot\rho|_{H}$. Then i) implies that there is a finite index subgroup $G^{\gamma}\subset G$ containing $H$ such that $\rho|_{G^{\gamma}}=\gamma\cdot\rho|_{G^{\gamma}}$. We define $G^{\prime}=\cap_{\gamma\in{\rm Gal}(\tilde{K}/K)}G^{\gamma}$ which contains $H$ and is a subgroup of finite index in $G$. Then $\rho|_{G^{\prime}}$ is ${\rm Gal}(\overline{K}/K)$ invariant thus descends to $K$.
@@
Assume that $\widehat{H}\to\widehat{G}$ is injective. Then finite direct sums of algebraically isomonodromic representations with semi-simple algebraic monodromy are algebraically isomonodromic. Direct summands of algebraically isomonodromic representations with semi-simple algebraic monodromy are algebraically isomonodromic.
@@
The second part is easy. For the first part we can assume without loss of generality that $K$ is algebraically closed and we can argue as in the proof of Theorem 4.1 reducing to the observation that direct sums of admissible representations relative to $H$ are admissible relative to $H$. ∎
@@ 5.1. Polarized variations of ${\mathbb{C}}$-Hodge structure
and the induced filtration is a -${\mathbb{C}}\text{-}{\sf PHS}$
@@
A polarized variation of ${\mathbb{C}}$-Hodge structure (-${\mathbb{C}}\text{-}{\sf PVHS}$) on a complex manifold $X$ is given by a triple $(\mathbb{L},\mathsf{F},Q)$, where $\mathbb{L}$ is a complex local system on $X$, $\mathsf{F}\subset\mathbb{L}\otimes_{\mathbb{C}}\mathcal{{\mathcal{O}}}_{X}$ is a finite filtration by holomorphic subbundles and $Q\colon\mathbb{L}\times\overline{\mathbb{L}}\to{\mathbb{C}}$ is a Hermitian perfect pairing, where we assume that the Hodge-Riemann relations at each point of $X$ are statisfied and Griffiths transversality $\nabla(\mathsf{F}^{a})\subset\Omega^{1}_{X}(\mathsf{F}^{a-1})$ holds.
@@
    if $\mathbb{L}$ underlies a $\mathbb{Z}$-local system then the algebraic monodromy group $\mathrm{Mo}(\mathbb{L},x)^{0}$ is semi-simple and the monodromy at infinity of $\mathbb{L}$ is quasi-unipotent.
@@
Let $f\colon Y\to X$ be either a finite étale covering or a Zariski open embeding (both with dense image) of quasi-projective complex manifolds. Let $\mathbb{L}$ be a ${\mathbb{C}}$-local system on $X$. Then $\mathbb{L}$ underlies a -${\mathbb{C}}\text{-}{\sf PVHS}$ if and only if $f^{*}\mathbb{L}$ underlies a -${\mathbb{C}}\text{-}{\sf PVHS}$.
@@
The case of a Zariski open embeding follows from the nilpotent orbit theorem [CK89, (2.1)]. ∎
@@
    for $0<|z|\ll 1$ the filtration $\mathsf{F}_{z}$ is a pure Hodge structure polarized by $Q_{z}$;
@@
    $\mathsf{F}$ extends to a filtration by holomorphic subbundles of $\mathcal{E}$ and its fibre over 00 induces a mixed Hodge structure polarized by $N$ and $Q_{0}$, i.e. a nilpotent orbit by [CK89, (2.3)].
@@
There is a version of Proposition 5.3 with continuous dependence on parameters. We later need 2) $\Rightarrow$ 1) with parameters, i.e. if the family of holomorphic filtrations $(\mathsf{F}(s))_{s\in V}$ of $\mathcal{E}$ depends continuously on parameters $s\in V$ with $V\subset\mathbb{R}^{m}$ open, $0\in V$, and for every $s\in V$ the filtration $\mathsf{F}_{0}(s)$ is a mixed Hodge structure on $\mathcal{E}_{0}$ polarized by $N$ and $Q_{0}$, then there exists $\epsilon>0$ and a neighborhood $U\subset V$ of 00 such $\mathsf{F}_{z}(s)$ is a pure Hodge structure polarized by $Q_{z}$ for $0<|z|<\epsilon$ and $s\in U$, see the argument in [CK89, Sec. 4] which extends immediately to the parametrized situation.
@@
Note that the classical literature, see [CK89], the degeneration of Hodge structure, and as a consequence Deligne’s semi-simplicity theorem, are studied for -${\mathbb{R}}\text{-}{\sf PVHS}$ with quasi-unipotent monodromy at infinity. Deligne remarked that the theory generalizes to arbitrary -${\mathbb{C}}\text{-}{\sf PVHS}$ [Del87, 1.11]. Recently, this was documented in the one variable case in [SS22].
@@
Let $\overline{X}$ be a projective complex manifolds (with fixed ample line bundle). Let $D\subset\overline{X}$ be a simple normal crossings divisor with complement $X$. We recall following Simpson [Sim92, Lem. 4.1] and T. Mochizuki [Moc06, 10.1] a citerion for a simple complex local system $\mathbb{L}$ on $X$ to underly a -${\mathbb{C}}\text{-}{\sf PVHS}$. Let $E$ be the $\mathcal{C}^{\infty}$-bundle underlying $\mathbb{L}$ with its flat connection $\nabla$.
[…]

### v5

@@ 4.2. Admissible representations
In this subsection we assume that $K$ is algebraically closed of characteristic zero. A representation $\rho_{H}\colon H\to{\mathrm{GL}}(V)$ is called admissible if it is semisimple and all simple constituents of $\rho_{H}$ have finite determinant. A representation $\rho\colon G\to{\mathrm{GL}}(V)$ is called admissible relative to $H$ (or an admissible extension of $\rho|_{H}$) if $\rho$ is admissible and all its simple constituents stay simple when restricted to $H$.
@@
The algebraic monodromy group $\mathrm{Mo}(\rho)^{0}$ of $\rho$ is semisimple if and only if $\rho|_{G^{\prime}}$ is admissible for all subgroups of finite index $G^{\prime}\subset G$.
@@
Assume the algebraic monodromy group $\mathrm{Mo}(\rho)^{0}$ is semisimple. Then for all finite index subgroups $G^{\prime}\subset G$, $\rho(G^{\prime})\subset\rho(G)$ has finite index, thus $\mathrm{Mo}(\rho|_{G^{\prime}})\subset\mathrm{Mo}(\rho)$ has finite index as well. This implies the equality $\mathrm{Mo}(\rho|_{G^{\prime}})^{0}=\mathrm{Mo}(\rho)^{0}$.
In particular, $\mathrm{Mo}(\rho|_{G^{\prime}})^{0}$ is reductive, thus by Weyl’s theorem, $\rho|_{G^{\prime}}$ as a representation of the algebraic group ${\rm Mo}(\rho|_{G^{\prime}})^{0}$ is semisimple, thus $\rho|_{G^{\prime}}$ itself is semisimple. As ${\rm Mo}(\rho)^{0}=[{\rm Mo}(\rho)^{0},{\rm Mo}(\rho)^{0}]$, all rank one algebraic representations of ${\rm Mo}(\rho|_{G^{\prime}})$ are finite. Thus, $\rho|_{G^{\prime}}$ is admissible.
Assume conversely that $\rho|_{G^{\prime}}$ is admissible for all subgroups of finite index $G^{\prime}\subset G$. Upon replacing $G$ by the preimage of ${\rm Mo}(\rho)^{0}(K)$, which is a finite index subgroup in $G$ (recall $K$ is algebraically closed), we may assume that $\mathrm{Mo}(\rho)$ is connected. As the unipotent radical $U$ of $\mathrm{Mo}(\rho)$ is normal, by Clifford’s theorem $\rho|_{U}$ is still semisimple. So by Engel’s theorem, we conclude that $U$ is trivial. The radical $R$ of $\mathrm{Mo}(\rho)$ is then a central torus, so by Schur’s lemma it acts by a character $\lambda\colon R\to\mathbb{G}_{m}$ on each simple $G$-subrepresentation $V_{i}\subset V$. As the action on $\det(V_{i})$ is finite, a power of $\lambda$ is trivial, so $\lambda$ is trivial. As the action of $R$ on $V$ is faithful, $R$ is trivial and ${\rm Mo}(\rho)^{0}$ is semisimple. ∎
@@
Finite direct sums of representations with semisimple algebraic monodromy groups have semisimple algebraic monodromy groups. Direct summands of representations with semisimple algebraic monodromy group have a semisimple algebraic monodromy group.
@@
By Lemma 4.2 it is sufficient to prove this for “semisimple algebraic monodromy group” replaced by “admissible in restriction to all finite index subgroups”, for which it is trivial. ∎
@@
for all $h\in H$ and $g\in G$, and as $\rho|_{H}$ is simple, Schur’s lemma implies that $\tau(g)\in K^{\times}$ for all $g\in G$. In fact $\tau(g)\in\mu_{N}(K)$ for some $N>0$ as the determinants of $\rho$ and $\tilde{\rho}$ are finite. This implies that $\tau\colon G\to\mu_{N}(K)\subset{\mathrm{GL}}(V)$ is a homomorphism. Thus, $\rho=\tilde{\rho}$ in restriction to the kernel $G^{\prime}\subset G$ of $\tau$, which is a finite index subgroup. ∎
@@ 4.3. Algebraically isomonodromic representations
If $\rho$ is algebraically isomonodromic relative to $H$, then $\mathrm{Mo}(\rho|_{H})\to\mathrm{Mo}(\rho)$ is essentially an isomorphism. In particular, $\mathrm{End}_{\rho}(V)\xrightarrow{\sim}\mathrm{End}_{\rho|_{H}}(V)$ is essentially an isomorphism.
@@
Assume $K$ is algebraically closed. If $\rho\colon G\to{\mathrm{GL}}(V)$ is algebraically isomonodromic relative to $H$ and $\mathrm{Mo}(\rho)^{0}$ semisimple, then $\rho$ is essentially admissible relative to $H$.
@@
Assume $K$ is algebraically closed. Let $\rho\colon G\to{\mathrm{GL}}(V)$ be a represenation such that $\mathrm{Mo}(\rho|_{H})$ is connected and semisimple. Then $\rho$ is essentially admissible relative to $H$ if and only if $\rho$ is algebraically isomonodromic relative to $H$.
@@
For $\Rightarrow$” we first observe that by replacing $G$ by the preimage of $\mathrm{Mo}(\rho)^{0}$, which is a subgroup of finite index containing $H$, we may assume that $\mathrm{Mo}(\rho)$ is connected. The argument in the proof of Lemma 4.2 shows that $\mathrm{Mo}(\rho)$ is semisimple. If the embedding $\mathrm{Mo}(\rho|_{H})\to\mathrm{Mo}(\rho)$ is not an isomorphism then there exists a non-trivial normal connected algebraic subgroup $W\subset\mathrm{Mo}(\rho)$ centralizing $\mathrm{Mo}(\rho|_{H})$ such that
@@
is surjective with finite kernel, see e.g. [Hum75, 14.2]. The $W$ action stabilizes every simple $\rho$-subrepresentation $V^{\prime}\subset V$ and commutes with the action of $H$ on $V^{\prime}$, which is simple. So by Schur’s lemma $W$ acts by a character on $V^{\prime}$, but all characters are trivial as $W$ is semisimple. So $W$ has to be trivial as it acts faithfully on $V$. ∎
@@
Assume that $K$ is algebraically closed. Let $\rho_{H}\colon H\to{\mathrm{GL}}(V)$ be a representation with semisimple algebraic monodromy group. Let $G^{\prime}\subset G$ be a subgroup of finite index, $H^{\prime}=H\cap G^{\prime}$. If $\mathrm{Mo}(\rho|_{H^{\prime}})$ is connected and $\rho|_{H^{\prime}}$ has an admissible extension $\rho^{\prime}\colon G^{\prime}\to{\mathrm{GL}}(V)$ relative to $H^{\prime}$, then $\rho$ essentially has an algebraically isomonodromic extension $\rho\colon G\to{\mathrm{GL}}(V)$ relative to $H$.
@@
After replacing $G^{\prime}$ by a subgroup of finite index we may assume that $G^{\prime}\subset G$ is normal, thus $H^{\prime}\subset H$ is normal as well. Replace $G$ by the preimage of $H/H^{\prime}\subset G/G^{\prime}$ via $G\to G/G^{\prime}$. Then by definition $G:G^{\prime}]=[H:H^{\prime}]$. We set
@@
We show in the sequel that there is a direct summand of $\tilde{\rho}$ which extends $\rho_{H}$. First observe that $\tilde{\rho}$ is algebraically isomonodromic relative to $H$. Indeed, it suffices to check that $\tilde{\rho}|_{G^{\prime}}$ is algebraically isomonodromic relative to $H^{\prime}$. Noting that $\tilde{\rho}|_{H^{\prime}}$ is a direct sum of copies of $\rho_{H}|_{H^{\prime}}$, so $\mathrm{Mo}(\tilde{\rho}|_{H^{\prime}})$ is connected and semisimple, by Proposition 4.10 we have to show that $\tilde{\rho}|_{G^{\prime}}$ is admissible relative to $H^{\prime}$. This holds as $\tilde{\rho}|_{G^{\prime}}$ it is a direct sum of $H$-conjugates of $\rho^{\prime}$ and $\rho^{\prime}$ is admissible relative to $H^{\prime}$. On the other hand,
@@
[…]
@@
Assume that $\widehat{H}\to\widehat{G}$ is injective. Then finite direct sums of algebraically isomonodromic representations with semisimple algebraic monodromy are algebraically isomonodromic. Direct summands of algebraically isomonodromic representations with semisimple algebraic monodromy are algebraically isomonodromic.
@@
The second part is easy. For the first part we can assume without loss of generality that $K$ is algebraically closed, and we can argue as in the proof of Theorem 4.1 reducing to the observation that direct sums of admissible representations relative to $H$ are admissible relative to $H$. ∎
@@ 5.1. Polarized variations of $\mathbb{C}$}-Hodge structure
and the induced filtration is a -${\mathbb{C}}\text{-}{\sf PHS}$ .
@@
A polarized variation of $\mathbb{C}$}-Hodge structure (-${\mathbb{C}}\text{-}{\sf PVHS}$) on a complex manifold $X$ is given by a triple $(\mathbb{L},\mathsf{F},Q)$, where $\mathbb{L}$ is a complex local system on $X$, $\mathsf{F}\subset\mathbb{L}\otimes_{\mathbb{C}}\mathcal{{\mathcal{O}}}_{X}$ is a finite filtration by holomorphic subbundles and $Q\colon\mathbb{L}\times\overline{\mathbb{L}}\to{\mathbb{C}}$ is a hermitian perfect pairing, where we assume that the Hodge-Riemann relations at each point of $X$ are satisfied and Griffiths transversality $\nabla(\mathsf{F}^{a})\subset\Omega^{1}_{X}(\mathsf{F}^{a-1})$ holds.
@@
    if $\mathbb{L}$ underlies a $\mathbb{Z}$-local system then the algebraic monodromy group $\mathrm{Mo}(\mathbb{L},x)^{0}$ is semisimple and the monodromy at infinity of $\mathbb{L}$ is quasi-unipotent.
@@
Let $f\colon Y\to X$ be either a finite étale covering or a Zariski open embedding (both with dense image) of quasi-projective complex manifolds. Let $\mathbb{L}$ be a $\mathbb{C}$}-local system on $X$. Then $\mathbb{L}$ underlies a -${\mathbb{C}}\text{-}{\sf PVHS}$ if and only if $f^{*}\mathbb{L}$ underlies a -${\mathbb{C}}\text{-}{\sf PVHS}$.
@@
The case of a Zariski open embedding follows from the nilpotent orbit theorem [CK89, (2.1)]. ∎
@@
    for $0<|z|\ll 1$, $(\mathcal{E}_{z},{\mathsf{F}}_{z},Q_{z}$) is a pure polarized Hodge structure;
@@
    $\mathsf{F}$ extends to a filtration by holomorphic subbundles of $\mathcal{E}$ and $(\mathcal{E}_{0},{\mathsf{F}}_{0},N,Q_{0})$ is a polarized mixed Hodge structure, i.e. it is a nilpotent orbit by [CK89, (2.3)].
@@
There is a version of Proposition 5.3 with continuous dependence on parameters. We later need 2) $\Rightarrow$ 1) with parameters, i.e. if the family of holomorphic filtrations $(\mathsf{F}(s))_{s\in V}$ of $\mathcal{E}$ depends continuously on parameters $s\in V$ with $V\subset\mathbb{R}^{m}$ open, $0\in V$, and for every $s\in V$ the filtration $\mathsf{F}_{0}(s)$ is a mixed Hodge structure on $\mathcal{E}_{0}$ polarized by $N$ and $Q_{0}$, then there exists $\epsilon>0$ and a neighborhood $U\subset V$ of 0 such that $(\mathcal{E}(s),\mathsf{F}_{z}(s))$ is a pure Hodge structure polarized by $Q_{z}$ for $0<|z|<\epsilon$ and $s\in U$, see the argument in [CK89, Sec. 4] which extends immediately to the parametrized situation.
@@
Note that in the classical literature, see [CK89], the degeneration of Hodge structure, and as a consequence Deligne’s semisimplicity theorem, are studied for -${\mathbb{R}}\text{-}{\sf PVHS}$ with quasi-unipotent monodromy at infinity. Deligne remarked that the theory generalizes to arbitrary -${\mathbb{C}}\text{-}{\sf PVHS}$ [Del87, 1.11]. Recently, this was documented in the one variable case in [SS22].
@@
Let $\overline{X}$ be a projective complex manifold (with fixed ample line bundle). Let $D\subset\overline{X}$ be a simple normal crossings divisor with complement $X$. We recall following Simpson [Sim92, Lem. 4.1] and T. Mochizuki [Moc06, 10.1] a citerion for a simple complex local system $\mathbb{L}$ on $X$ to underly a -${\mathbb{C}}\text{-}{\sf PVHS}$. Let $E$ be the $\mathcal{C}^{\infty}$-bundle underlying $\mathbb{L}$ with its flat connection $\nabla$.
T. Mochizuki and Jost–Zuo construct a tame, purely imaginary pluri-harmonic metric $h$ on $E$ which is unique up to a factor in $\mathbb{R}_{>0}$, see [Moc07, Part 5] and [Moc09]. This allows us to write $\nabla=\partial_{E}+\overline{\partial}_{E}+\theta+\theta^{\dagger}$. Here $\theta$ is a $(1,0)$-form with values in $\mathit{End}(E)$ with adjoint $\theta^{\dagger}$ with respect to $h$ and $\partial_{E}+\overline{\partial}_{E}$ is a connection compatible with $h$. The associated Higgs bundle $(E,\overline{\partial}_{E},\theta)$ has a canonical parabolic structure at infinity such that $\theta$ has only simple poles and purely imaginary eigenvalues of its residues.